OpenStudy (linda):

sigma (E) with 3 on top. n=0 on bottom. what does this mean??

OpenStudy (linda):

on the side on sigma, it is sin\[\pi\] over 4

OpenStudy (anonymous):

\[\sum_{n=0}^{3}\sin \pi/4\] ?

OpenStudy (anonymous):

Look something like that?

OpenStudy (linda):

theres an n before the pi symbol, but other than that its right

OpenStudy (anonymous):

\[Ok so \sum_{n=0}^{3}\sin (n*\pi/4)\] So like this?

OpenStudy (anonymous):

Well without the okso at the beginning.

OpenStudy (linda):

yeah thats right

OpenStudy (anonymous):

The sigma represents a summation. So if it were \[\sum_{n=0}^{3}n\] we start with n=0 plug that value in and keep adding until we reach the limit which is the top value on the sigma, 3.

OpenStudy (anonymous):

So we'd have 0+1+2+3 and then stop.

OpenStudy (linda):

so when n=0, the whole thing would equal 0 right?

OpenStudy (anonymous):

Yes. You start with whatever n equals under the sigma and plug it into the function to the left of the sigma and then go to the next sequential n value and add it to that result.

OpenStudy (linda):

so do i just start with n=0 and I get 0, then n=1 I plug it into n*pi over 4 and that equals .785 and so I add that to the zero.... then i go on to n=2 and keep adding on to the .785......

OpenStudy (anonymous):

exactly!

OpenStudy (linda):

great thanks!

OpenStudy (anonymous):

good method...

OpenStudy (anonymous):

Do u still have doubts @linda ?